OpenStudy (anonymous):

Choose the equation below whose axis of symmetry is x = 2. y = x2 + 4x + 2 y = x2 − 4 y = x2 − 2 y = x2 − 4x + 2

OpenStudy (mertsj):

Calculate -b/2a until you find one where the result is 2

OpenStudy (anonymous):

The formula for the axis of symmetry is x = -b/2a. Use y = ax² + bx + c, along with the first-mentioned formula to determine the correct equation. 1) y = x² + 4x + 2 a = 1 b = 4 c = 2 x = -4/(2) ==> -2 [Not what we want!] 2. y = x² - 4 a = 1 b = 0 c = -4 x = -0/2 ==> 0 [This doesn't match with x = 2!] 3. Same as equation #2. The axis of symmetry is x = 0. 4. y = x² - 4x + 2 a = 1 b = -4 c = 2 x = -(-4)/(2) ==> 4/2 ==> 2 [Yes! That's the equation with the axis of symmetry of x = 2!]

OpenStudy (anonymous):

You get it?